The core of the optical fiber is glass with a refractive index of 1.65. The fiber sheath is also glass, but its refractive index is 1.50.  At what angle x with respect to the central axis of the fiber must the laser beam strike the end of the fiber core in order for it to be totally reflected at the core-sheath interface? The cutting surface of the fiber end is perpendicular to the axis. 



1
Expert's answer
2021-08-18T07:02:11-0400

The angle of total internal reflection

"\theta=\sin^{-1}\frac{1.50}{1.65}=65.4^{\circ}"